8-pack control system working group meeting

18 February 2002

In attendance: Chestnut, Hall, Kim, Nelson, Rock, S Smith

Notes by: J Nelson


(This look a lot like last week’s, but the contents are different.  I promise.  Unchanged contents are prefaced with (no change).)



-(no change)Progress from Kukhee: Critical work is done, a little minor work is left to do.

-need to do checkout with real signals

-need to check archiving of PMIN, PMAX, etc variables after we generate another to-be-archived list

-need from Andrew/Steve: of LD1 and LD2, which is closer to the enclosure (in the ADC channel list spreadsheet)



-The database and panel are done for the attenuation and pulse width control.  Small changes are needed as the DAC channels have changed.  Judy made the panel. 



DAC/ADC Issues

All of these issues are for later:


-Janice gave panels to look at to Ron for which phase PVs we care about.  He may work on this sooner.  TBD.  Previously it was decided that one could input any number but then it’d be given the equivalent value between +/- 180.

-TBD: The trickle down of the SLED variables.  This was specified at a previous meeting (numbers on SLED panel need to trickle down to I/Q and A/P PVs).  For processing, we might want this feature functional sooner rather than later. 

-TBD: the trickle down of the variables on the power distribution panel as well.  There are 4 PVs there called Klystron Sled-in phases.  These need to trickle down to the phase DACs as well.  This phase is to be added to the channel phase angle (IQ calib panel) and to the timestep phase before going out to the DAC.



-Done: Andrew will email Ron about the three bits for the multiplexer to find out which is the least significant bit.  The database is done.

-This is still to be done:  SWE will make a PV for a temporary DIO channel as a reset to Joe Frisch’s kludge RE Interlock.  They will send the name to Janice who will put a button on the panel with the pulse width and the klystron attenuation.  Its DIO channel assignment should be on Andrew’s previously sent list.  Also need a channel for status when it trips off.

These buttons should be on the pulsewidth/atten panel.

-(no change) The following item was deferred for a few weeks:

LLRF + NLCTA Ops need to sort out what the LLRF enable and the mod off buttons do


SCP Configs

-dbinstall Thursday – this will change all trigs to trbrs.  Janice needs to make mini-dbedit files for the trbrs will appropriate rep rates.



-Kukhee reported that all the waveform records and the fanouts have been created.

-The list of waveforms to archived was sent to Bob from Janice with Judy’s help.

-(no change) We deferred the following: longer term issue: what to do if reading out an SIS when a fault occurs?

-(no change) Right now the SISs are paused and read out with a user request only.  A read when a fault happens will also be added to this.



- Done/ongoing: Bob reported that the 8-pack variables are being archived and the archiver browser gets its data from the oracle database.  Over the next x weeks more  PVs will be added. 

-Bob will setup the browser to use the selection criteria for helping folks add channels.  Will be finished in about 2 weeks.

-Done: Janice will check the 8-pack main panel to make sure it uses cspaarch and ditto the SCP.

-Janice will also check the archiving of the waveforms and perhaps with Steve/Andrew’s help send in a fake signal.  To be done when Janice gets back.



-DONE: Janice will send a list of waveforms to be archived to Bob et al (see first item until SISs above)

-Bob is working on Matlab functions to allow users to access the archived waveforms.  See previous notes for specs.  Coming soon…

-(no change) SWE plans to implement the following two features requested by LLRF (Joe’s email):

have waveforms archived on (1) every trip and (2) when an acquire is done by a user

-(no change) Later SWE will work with Joe on his other two requests: archive every few minutes and when a change occurs in a selected set of values in EPICS (a pseudo-trip if you will)



-Done and checked: To be done this week: Kukhee will swap the 0-15 assignment (again).  Steve reported that the IQA channels are correct but the time steps are backwards. 



-All of the following will get going in the next couple weeks:

-Each IQA channel has 6 or 7 calibration numbers that are module dependent.  Gain/offset I,Q, quadrature angle and overall phase.  These numbers are used to change the SIS waveforms from I/Q to A/P.  These numbers will go on the IQA module panel. 

-Steve will email the equations for this calculation - soonish

-Kukhee needs to change the SIS converter subrecord and new PVs need to be created.

-Once names are decided upon, Janice can make the panel

-These 6 (or 7) numbers x 8 channels x 4 IQAs need to be archived at least once every 10 minutes.  The panel interface could just be one table for each IQA with 8 rows for the 8 channels and 6 columns with the calibration variables.


Processing algorithm

-(no change) Ron and Kukhee reported that the work is ongoing, fast loop should be done n the next 2 weeks, the slow processing (run-up) loop will come a few weeks after that.  First phase will be done in about 3 weeks – to be assessed.  The
”shut off button” will set pw to min and attn to 0. 

The LLRF atten/pw display will have “Joe Signal” and a “Joe Reset” buttons.


Peak Power Meter (PPM)

-The list of PV names has been decided. 

-Done and checked: Judy will make the database for the descriptive string.  She and Ron reported how to record a change in this string has been solved.  The information will be in error log.  39 chars will be sufficient.

-(no change) Tonee will tweak the labview program to send the waveform to epics once we have names – probably late this week

-(no change) Tonee and Janice will set up the necessary hardware in ESB (Enet box, cables, etc)

-Done: Judy will make new panels once names have been decided.  Waveform plot buttons were checked, and appear to be working.  Will test for sure when labview is ready.



-(no change) We’re currently waiting for Dick’s setup to settle down then we’ll do one more iteration of getting list, changing database, changing panels, programming the dcm, changing alarm handler and archiver stuff as needed



-Done: print button to Steve’s printer ok?

-Add Steve’s color printer to EPICS print list.  Soon, hopefully (needs printer subnet setup).

-(no change)Judy reported the crate profiles on the web are updated and look good.  Janice requested that channel assignment information also be available via the web, even it it’s just a link to the template file.  Judy and Ron will look into this.

-Andrew will tell Kukhee when he takes out the DIO module to modify it for the deadman.  no change: Janice requested that a bypass be incorporated into the design.  Is this needed?  Janice will ask Marc. 

-(no change) we deferred the following:

-public/private matlab issues to be resolved at a later date

-LLRF needs to decide what the reset button does.



Sort task list by person


Ron (and otherwise unassigned SWE tasks):

-PVs/database for 5 dac channels (pulse width, attenuation)

-PV/database for DIO reset of Joe Frisch’s kludge RE Interlock

-phase wraparound

-trickle down of sled variables

-trickle down of phase variables

-check status of 8-pack PVs being archived and the browser chooser.

-check status of switchover to Oracle archiver database

-database/PVs for 6 channels of calibration numbers for each IQA channel (also connect to Kukhee’s SIS converter subroutine)

-processing algorithm

-PVs/database for PPM descriptive string variable

-save of descriptive name for PPM channels



-swap VSTEP channel assignments (again)

-ADC work on going



-SCP timing config, TRBRs or TRIGs (make all the same!)

-checkout of waveform archiving

-check 8-pack main panel and scp panel for correct archive browser call (cspaarch)

-once names are set, make panel with 5 dac channels (pulse width and attenuation) and the DIO reset of Joe Frisch’s kludge RE Interlock

-panel for 6 calib numbers for each IQA channel

-PPM panel, work with Tonee on PPM labview and hardware



-check into making module channel assignments available on the web as well as the crate profiles




-pass on equation for the 6 calibration numbers for each IQA channel

-kludge interlock reset hardware

-email Ron about the DIO bit order for the IQA multiplexer



-matlab functions to fetch archived waveforms.